OpenStudy (anonymous):

What is the quotient: (3x2 + 8x – 3) ÷ (x + 3) ? Answer 3x + 5, r = 1 3x – 1 3x + 1 3x – 5

OpenStudy (anonymous):

Factor the numerator by grouping 3x^2+8x-3 =3x^2+9x-x-3 =3x(x+3)-(x+3) =(x+3)(3x-1)

OpenStudy (anonymous):

Use the long division method to find the answer :) \[3x ^{2}+8x-3/x+3\] You get, 3x-1 as the quotient. :)

OpenStudy (anonymous):

thanks yall. :)